OpenStudy (anonymous):

WILL GIVE MEDAL!!!!! An archer pulls a bowstring back 0.440 m, and the spring constant of the bow is 545 N/m. a.) Find the amount of work the archer did to draw the bow. b.) Find the speed an arrow of mass 0.060 kg will have when it is shot from the bow. Assume no losses from friction.

OpenStudy (anonymous):

@oksuz_ can you help?

OpenStudy (anonymous):

W=Ep=(kx^2)/2 Ep=Ek= (mv^2)/2 use these...

OpenStudy (anonymous):

I don't know K or x so I can't solve

OpenStudy (anonymous):

they are given.. K= 545 N/m X= 0.440 m

OpenStudy (anonymous):

so work=545(.440^2) = 105.5 J

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

Sorry, I stink at physics.